OpenStudy (goformit100):

The number of boys and girls in a class are in the ratio 7:5. The number of boys is 8 more than the number of girls. What is the total class strength ?

OpenStudy (goformit100):

@AkashdeepDeb

OpenStudy (anonymous):

Let Boys = b Girls = g The number of boys and girls in a class are in the ratio 7:5 b:g=7:5 b/g=7/5 b5=g7 b5/5=g7/5 b=g7/5...............(1) the number of boys is 8 more than the number of girls b=g+8.............(2) Put the value of b from (1) to (2) g7/5=g+8 5*(g7/5)=5*(g+8) g7=5g+40 g7-g5=40 2g=40 2g/2=40/2 g=20 Put the value of g in (2) b=20+8 b=28 Boys are 28 and Girls are 20

OpenStudy (anonymous):

As said in the bottom Boys are 28 and Girls are 20

OpenStudy (anonymous):

If this helped please close the question.

OpenStudy (goformit100):

is it in total 48 students ?

OpenStudy (anonymous):

28+20= 48 soo yea

OpenStudy (goformit100):

Thanks

OpenStudy (anonymous):

NP.